# Clinical-R1: Empowering Large Language Models for Faithful and Comprehensive Reasoning with Clinical Objective Relative Policy Optimization

### v1

@@ 3 Methodology
Current post-training methods such as GRPO optimize only for correctness, without constraining how reasoning is formed. As a result, models often produce fragmented or shortcut explanations. This makes the reasoning process unreliable and hard to validate, especially in medicine, where faithfulness and completeness are as critical as accuracy. Clinicians, in contrast, rely on Dual Process Thinking (djulbegovic2012dual): an analytical process (System 2) that systematically evaluates hypotheses, and an intuitive process (System 1) that draws on experience and pattern recognition. Physicians often use System 1 to generate hypotheses and System 2 to verify or refine them, ensuring that intuitive insights remain grounded in analytical reasoning. To align LLMs with this cognitive structure, as shown in Table LABEL:tab:arch, we introduce the Clinical Reasoning Reward, which enforces a structured reasoning format reflecting these two systems. The model must separate analytical reasoning and conclusion using <dx>, </dx>, <conclusion>, and </conclusion>tags. The <dx>section captures hypothesis-driven reasoning, while the <conclusion>section synthesizes these insights into a conclusion. Additional reward is given when the conclusion explicitly references analytical elements from the dx section, promoting faithfulness and comprehensiveness.
@@
Table LABEL:tab:arch shows our designed rule-based reward that follows the convention created by GRPO with small variations. The reward is the summation of three parts: accuracy reward, clinical reasoning reward, and consistency reward.
@@
  • Clinical reasoning reward: We require the model output to generate reasoning and the answer between the <dx>, </dx>, <conclusion>, and </conclusion>tags. Within the dx tags, we explictly ask for a separation of an intuitive process (System 1) and an analytical process (System 2). If the response is in such format, a reward of $1$ is given. To further encourage faithfulness and comprehensiveness, if the <conclusion>part cross-references the <dx>part, an additional reward of $0.5$ is given. Otherwise, the clinical reasoning reward is 0.
@@ 4.2 Baseline and Evaluation Metrics
We compare our Qwen2.5-3B-based Clinical-R1 with the original pretrained model with CoT prompting. The CoT setting uses the same reasoning template as the CRPO sampling prompt shown in Table LABEL:tab:arch. This baseline establishes the performance of the base model before any distillation or reinforcement learning.
@@ 5.1 Results
As shown in Figure 3, 4, across three benchmarks (MedQA in-domain; MedMCQA and MedXpertQA out-of-domain), CRPO-trained models match or surpass GRPO in answer accuracy. The cold-start + CRPO variant (Clinical-R1-3B) yielded the strongest competitive results. Majority voting adds little in already low-variance settings, suggesting that CRPO promotes self-consistent generations where ensembling offers limited marginal benefit (Table 1). The results show the effectiveness of our approach.

### v2

@@ 3 Methodology
Current post-training methods such as GRPO optimize only for correctness, without constraining how reasoning is formed. As a result, models often produce fragmented or shortcut explanations. This makes the reasoning process unreliable and hard to validate, especially in medicine, where faithfulness and completeness are as critical as accuracy. Clinicians, in contrast, rely on Dual Process Thinking (djulbegovic2012dual): an analytical process (System 2) that systematically evaluates hypotheses, and an intuitive process (System 1) that draws on experience and pattern recognition. Physicians often use System 1 to generate hypotheses and System 2 to verify or refine them, ensuring that intuitive insights remain grounded in analytical reasoning. To align LLMs with this cognitive structure, as shown in Figure 2, we introduce the Clinical Reasoning Reward, which enforces a structured reasoning format reflecting these two systems. The model must separate analytical reasoning and conclusion using <dx>, </dx>, <conclusion>, and </conclusion>tags. The <dx>section captures hypothesis-driven reasoning, while the <conclusion>section synthesizes these insights into a conclusion. Additional reward is given when the conclusion explicitly references analytical elements from the dx section, promoting faithfulness and comprehensiveness.
@@
Figure 2 shows our designed rule-based reward that follows the convention created by GRPO with small variations. The reward is the summation of three parts: accuracy reward, clinical reasoning reward, and consistency reward.
@@
  • Clinical reasoning reward: We require the model output to generate reasoning and the answer between the <dx>, </dx>, <conclusion>, and </conclusion>tags. Within the <dx>tags, we explicitly ask for a separation of an intuitive process (System 1) and an analytical process (System 2). If the response is in such format, a reward of $1$ is given. To further encourage faithfulness and comprehensiveness, if the <conclusion>part cross-references the <dx>part, an additional reward of $0.5$ is given. Otherwise, the clinical reasoning reward is 0.
@@ 4.2 Baseline and Evaluation Metrics
We compare our Qwen2.5-3B-based Clinical-R1 with the original pretrained model with CoT prompting. The CoT setting uses the same reasoning template as the CRPO sampling prompt shown in Figure 2. This baseline establishes the performance of the base model before any distillation or reinforcement learning.
@@ 5.1 Results
As shown in Figure 3 and 4, across three benchmarks (MedQA in-domain; MedMCQA and MedXpertQA out-of-domain), CRPO-trained models match or surpass GRPO in answer accuracy. The cold-start + CRPO variant (Clinical-R1-3B) yielded the strongest competitive results. Majority voting adds little in already low-variance settings, suggesting that CRPO promotes self-consistent generations where ensembling offers limited marginal benefit (Table 1). The results show the effectiveness of our approach.
@@ 6 Conclusion and Limitation
Acknowledgements
DAC was funded by an NIHR Research Professorship; a Royal Academy of Engineering Research Chair; and the InnoHK Hong Kong Centre for Cerebro-cardiovascular Engineering (COCHE); and was supported by the National Institute for Health Research (NIHR) Oxford Biomedical Research Centre (BRC) and the Pandemic Sciences Institute at the University of Oxford. The Applied Digital Health (ADH) group at the Nuffield Department of Primary Care Health Sciences is supported by the National Institute for Health and Care Research (NIHR) Applied Research Collaboration Oxford and Thames Valley at Oxford Health NHS Foundation Trust. The views expressed are those of the author(s) and not necessarily those of the NHS, the NIHR or the Department of Health and Social Care. FL was funded by the Clarendon Fund and the Magdalen Graduate Scholarship. HZ was funded by the Clarendon Fund, the Department of Engineering Science Studentship, and the Frederick Brodckhues Scholarship. BMS is funded by the Rhodes Trust under the Rhodes scholarship. We sincerely thank all the editors and reviewers for their constructive comments and suggestions that substantially improved this paper.